# Exploring Nonreciprocal Noise Transfer under Onsager-Casimir Symmetry in Synthetic-Field Optomechanics

### v1

@@ V Conclusion
In summary, we have conducted a theoretical investigation of a three-mode closed-loop optomechanical system that exhibits a synthetic magnetic field. While the Onsager-Casimir reciprocity relations remain valid, the introduction of a loop phase effectively breaks time-reversal symmetry, thereby enabling non-reciprocal propagation of backaction noise within the system. By analyzing the PSD of the internal fluctuations, we demonstrated that this non-reciprocity is manifested as a directional transfer of noise between the mechanical resonators. This directional behavior is a direct consequence of the phase-dependent interference in the loop structure, which redistributes correlated noise asymmetrically across the system’s resonators.
@@
Importantly, this non-reciprocal noise transfer can be exploited for practical applications, such as enhancing the performance of optomechanical sensors. In particular, increasing the intermechanical coupling constant leads to improved sensitivity by amplifying the noise contrast between the resonators. Conversely, if the primary goal is to maximize non-reciprocity rather than sensitivity, a smaller intermechanical coupling becomes advantageous, as it enhances the asymmetry in the noise flow. These findings highlight the tunability of non-reciprocal effects in such systems and offer valuable insights for the design of next-generation quantum sensors and signal processing devices.
@@ Acknowledgements.
We thank İnanç Adagideli for fruitful discussions. B.S.E. and C.B. acknowledge support from Air Force Office of Scientific Research (AFOSR) Grant No. FA9550-22-1-0444. S.K.O. acknowledges support from AFOSR Multidisciplinary University Research Initiative (MURI) Award No. FA9550-21-1-$02$.
APPENDIX A: RESULTS UNDER DIFFERENT INTERMECHANICAL COUPLINGS
@@
In this Appendix, we elaborate on the effects of proximity to either of the EPs by varying the intermechanical coupling constant. For this purpose, we replot some of the figures in the Results section for different $\mu$ values. In Fig. A.1, we plot intracavity fluctuations and the noise contribution of the second mechanical resonator to cavity-output PSD at $\mu_{EP,1}$ to compare with Fig. 2. In comparison, the double peak characteristic in PSD when $\phi_{\ell}=\pi/2,3\pi/2$ is not fully developed at $\mu_{EP,1}$, whereas the noise power increases slightly.
@@
Figure A.1: (a) Optical intracavity PSD, (b) individual noise contribution of the second mechanical resonator to the cavity-output noise spectrum for three different loop-phases, $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$ at $\mu=\mu_{EP,1}$.
@@
In Fig. A.2, we examine internal mechanical fluctuations at the $\mu_{EP,1}$ compared to results at the $\mu_{EP,2}$. This outcome aligns well with the findings in Fig. 5, which shows enhanced sensitivity for higher values of the mechanical coupling constant.
Figure A.2: PSD of (a) first and (b) second mechanical resonators as a function of frequency for $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=\mu_{EP,1}$.
@@
In Figs. A.3-A.7, we replot Fig. 4 to analyze the differences depending on the proximity of EPs. These figures serve to visually support the discussion presented in Figs. 6 and 7. The non-reciprocity measure given in Eq.(7) is defined as the difference between the integrated PSD. Hence, it increases for small values of $\mu$ since noise power increases. However, for larger mechanical coupling constant values, noise power decreases, and as a result, integrated noise power also decreases. Another notable feature is the emergence of a double peak in the PSD as the system approaches the EPs, which can also be inferred from Fig. 7.
@@
Figure A.3: Noise transfer pathways between the two mechanical resonators. (a) Contribution of the second resonator’s fluctuations to the momentum-noise spectrum of the first resonator, (b) contribution of the first resonator’s fluctuations to the momentum-noise spectrum of the second resonator for $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=0.5\mu_{EP,1}$.
Figure A.4: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=\mu_{EP,1}$.
Figure A.5: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=1.5\mu_{EP,1}$.
Figure A.6: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=1.5\mu_{EP,2}$.
Figure A.7: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=2\mu_{EP,2}$.

### v2

@@ V Conclusions
In summary, we have conducted a theoretical investigation of a three-mode closed-loop optomechanical system that exhibits a synthetic magnetic field. While the Onsager-Casimir reciprocity relations remain valid, the introduction of a loop phase effectively breaks time-reversal symmetry, thereby enabling nonreciprocal propagation of backaction noise within the system. By analyzing the PSD of the internal fluctuations, we demonstrated that this nonreciprocity is manifested as a directional transfer of noise between the mechanical resonators. This directional behavior is a direct consequence of the phase-dependent interference, which redistributes correlated noise asymmetrically across the system’s resonators. An equally vital point is that the experimental realization of this theoretical proposal is currently feasible. For instance, the loop phase can be imparted as a phase difference between the two mechanical resonators by a pair of vibrating dielectric membranes embedded in a Fabry-Pérot cavity, with their individual radiative coupling amplitudes and phases tunable through the membranes’ absolute and relative positions [84, 85]. Furthermore, our scheme demands a strong intermechanical coupling constant, which has been demonstrated in earlier experiments as reviewed in the preceding experimental-realization Sec. III.
@@
Importantly, this nonreciprocal noise transfer can be exploited for practical applications, such as enhancing the performance of optomechanical sensors. In particular, increasing the intermechanical coupling constant leads to improved sensitivity by amplifying the noise contrast between the resonators. Conversely, if the primary goal is to maximize nonreciprocity rather than sensitivity, a smaller intermechanical coupling becomes advantageous, as it enhances the asymmetry in the noise flow. These findings highlight the tunability of nonreciprocal effects in such systems and offer valuable insights for the design of next-generation quantum sensors and signal-processing devices.
@@ Acknowledgements.
We thank İnanç Adagideli for fruitful discussions. B.S.E. and C.B. acknowledge support from Air Force Office of Scientific Research (AFOSR) Grant No. FA9550-22-1-0444. S.K.O. acknowledges support from Air Force Office of Scientific Research (AFOSR) Multidisciplinary University Research Initiative (MURI) Award No. FA9550-21-1-$02$.
Appendix A: Results under different mechanical coupling constants
@@
In this Appendix, we elaborate on the effects of proximity to either of the EPs by varying the intermechanical coupling constant. For this purpose, we replot some of the figures in the Results section for different $\mu$ values. In Fig. 8, we plot intracavity fluctuations and the noise contribution of the second mechanical resonator to cavity-output PSD at $\mu_{EP,1}$ to compare with Fig. 2. In comparison, the double peak characteristic in PSD when $\phi_{\ell}=\pi/2,3\pi/2$ is not fully developed at $\mu_{EP,1}$, whereas the noise power increases slightly.
@@
Figure 8: (a) Optical intracavity PSD, (b) individual noise contribution of the second mechanical resonator to the cavity-output noise spectrum for three different loop phases, $\phi_{\ell}=0,~\pi/2,~\pi$, and $3\pi/2$ at $\mu=\mu_{EP,1}$.
@@
In Fig. 9, we examine internal mechanical fluctuations at the $\mu_{EP,1}$ compared to results at the $\mu_{EP,2}$. This outcome aligns well with the findings in Fig. 5, which shows enhanced sensitivity for higher values of the mechanical coupling constant.
Figure 9: PSD of (a) first and (b) second mechanical resonators as a function of frequency for $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=\mu_{EP,1}$.
@@
In Figs. 10-14, we replot Fig. 4 to analyze the differences depending on the proximity of EPs. These figures serve to visually support the discussion presented in Figs. 6 and 7. The nonreciprocity measure given in Eq.(10) is defined as the difference between the integrated PSD. Hence, it increases for small values of $\mu$ since noise power increases. However, for larger mechanical coupling constant values, noise power decreases, and as a result, integrated noise power also decreases. Another notable feature is the emergence of a double peak in the PSD as the system approaches the EPs, which can also be inferred from Fig. 7.
@@
Figure 10: Noise transfer pathways between the two mechanical resonators. (a) Contribution of the second resonator’s fluctuations to the momentum-noise spectrum of the first resonator, (b) contribution of the first resonator’s fluctuations to the momentum-noise spectrum of the second resonator for $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=0.5\mu_{EP,1}$.
Figure 11: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=\mu_{EP,1}$.
Figure 12: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at $\mu=1.5\mu_{EP,1}$.
Figure 13: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=1.5\mu_{EP,2}$.
Figure 14: Noise transfer pathways between the two mechanical resonators $\phi_{\ell}=\pi/2$ and $3\pi/2$ at the $\mu=2\mu_{EP,2}$.
Appendix B: Root Loci of the upper half-plane eigenvalues
To identify the origin of asymmetry with respect to $\phi_{\ell}=\pi/2$ in Fig. 5, in this appendix we plot real and imaginary parts of the upper half-plane eigenvalues $(z(|\mu|,\phi_{\ell})=\alpha+i\omega,~\omega>0)$ of the drift matrix $\mathbf{M}$ as the loop phase varies $\phi_{\ell}\in[0,\pi]$ (for $\phi_{\ell}\in[\pi,2\pi]$ the same loci are retraced backwards). These are displayed at two different intermechanical coupling values chosen to be at either of the two EPs. For any $\phi_{\ell}$ there exist three eigenvalues corresponding to supermodes formed by the hybridization of two mechanical and one photonic modes. The conspicuous asymmetric behavior of these supermodes with respect to $\phi_{\ell}=\pi/2$ (marked with vertical dashed lines) in Fig.15 is the source of lack of this symmetry in the nonreciprocity measure in Fig. 5.
Figure 15: Real and imaginary parts of complex upper half plane eigenvalues $z(|\mu|,\phi_{\ell})$ of the drift matrix $\mathbf{M}$ as loop phase $\phi_{\ell}$ varies from 0 to $\pi$ at the EPs, $|\mu_{EP,1}|$ and $|\mu_{EP,2}|$. A vertical axis is placed at $\phi_{\ell}=\pi/2$ to highlight the asymmetry.